the perimeter of a rectangle is 108 cm. If the length is increased by 5 and the width is doubled. The perimeter becomes 182cm. Find it's dimension.
Answer by fractalier(6550)

Here perimeter P = 2L + 2W = 108 or L + W = 54 or L = 54 - W
Then in the second case,
2(L + 5) + 2(2W) = 182
Substitute the first equation into this one and get
2(54 - W + 5) + 4W = 182
2(59 - W) + 4W = 182
118 - 2W + 4W = 182
2W + 118 = 182
2W = 64
W = 32 cm
L = 54 - 32 = 22 cm